# UniGen: A Unified Framework for Textual Dataset Generation via Large Language Models

### v3

@@ 3.4 Post-Processing
Difficulty Enhancement. Given that the dataset is produced by LLMs, the complexity of the generated data is occasionally insufficient to challenge LLMs as their capabilities evolve. To address this, and inspired by prior research [14, 13], we implement several strategies to increase the data’s difficulty. These strategies are designed to elevate the challenges faced by LLMs in processing and responding to the data. The applied policies include: (1) Paraphrasing Question: Reformulate the phrasing to express the same idea with greater sophistication. (2) Adding Extra Context into Question: Integrate additional context or details that, while not directly aiding in the question’s resolution, enhance the question’s complexity. (3) Paraphrasing The Choices: Each option should be rephrased to reflect the same concept or idea as the original. The essence and meaning must be preserved. If an option cannot be paraphrased without altering its meaning, it should remain unchanged. (4) Adding A New Choice: Introduce a plausible but incorrect option to the existing choices to create ambiguity and require deeper understanding.
Group Checking. To mitigate the issue of high similarity among generated data items, a group-checking mechanism is implemented to identify and eliminate duplicates. Specifically, we utilize OpenAI’s text-embedding-ada-002 [59] to compute embeddings for all generated items. Let $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of generated data items, and $\mathbf{e}_{i}$ be the embedding of item $x_{i}$ computed via text-embedding-ada-002. We define the similarity matrix $\mathbf{S}$ where the element $s_{ij}$ is given by $s_{ij}=\sqrt{\sum_{k=1}^{d}(e_{ik}-e_{jk})^{2}}$, representing the Euclidean distance between the embeddings of items $x_{i}$ and $x_{j}$. Data items exhibiting a similarity exceeding a predefined threshold $\theta$ are filtered out to ensure diversity within the dataset. Formally, if $s_{ij}<\theta$ for any pair $(i,j)$, at least one of the items $x_{i}$ or $x_{j}$ is randomly removed from the final dataset.
4 Experiment and Application
@@ 4.1 Experimental Setup
To thoroughly evaluate the effectiveness of UniGen, we have carefully selected four representative benchmark datasets: GSM8K [19], TruthfulQA [72], MMLU [73], and HellaSwag [74]. Each dataset uniquely contributes to language model assessment, covering dimensions from mathematical problem-solving and factual accuracy verification to extensive language understanding and commonsense reasoning. We show the details of these four datasets in Appendix B. For dataset generation, we utilize GPT-4 [1], Claude3-Opus [2], and Llama3-70b [3], as these LLMs are among the most robust available, exhibiting exceptional ability to follow instructions. For benchmarking, our study utilizes eight cutting-edge and popular models from notable entities in the AI domain (the details are shown in Appendix B.), reflecting a mix of open-source and proprietary technologies. The number of generated data items and more details are shown in Appendix D. Note that difficulty enhancement is not applied to the generated data for benchmarking. We will discuss the effectiveness of difficult enhancement in subsection 4.3.
@@ 4.2 Characterizing Generated Data
Length. As depicted in 4(a), the length distribution of all generated datasets approximates a normal distribution. Notably, except for the HellaSwag dataset (as the length of the original HellaSwag dataset looks like a bimodal distribution), the distributions of other datasets closely resemble those of their respective original datasets. This similarity indicates that UniGen effectively mimics the distribution of the original data, thereby enhancing the reliability of the generated datasets.
Diversity. Analogous to the length distribution, the distribution of the self-BLEU score [75] (as depicted in 4(b))—a metric employed to assess text diversity—indicates that the diversity of the generated data closely aligns with that of the original dataset. This alignment underscores the exceptional capability of UniGen to replicate the diversity inherent in the original dataset, demonstrating its effectiveness in producing varied textual content. Additionally, we utilize the remote-clique metric, as applied in prior research [76], to measure the diversity of the generated data. The related statistics are presented in Table 2. Observations reveal that the remote-clique scores of the original and generated data are closely matched, with less than 10% variance, affirming that our generated data maintains a high level of diversity comparable to the original dataset.
Table 3: The knowledge richness comparison between different principles in DyVal 2 [13] and UniGen. The principle 1, 2, 3, and 4 are paraphrasing questions, paraphrasing choices, adding extra context to questions, and adding a new choice.
@@
UniGen 5.40% 3.30% 2.80% 3.83%
Knowledge Richness Introduced. In contrast to prior research [12, 13, 14], UniGen innovates by generating entirely new data items, rather than merely modifying existing answers. This approach introduces novel scenarios and knowledge. We assessed the knowledge richness of the data generated by UniGen and compared it to the previous study (i.e., Dyval2 [13]) by calculating the entity overlap rate—how many entities appear both in the generated and original data. A lower overlap rate indicates that the framework is introducing more new knowledge. According to our findings, presented in Table 3, UniGen demonstrates an average overlap rate of only 3.83%, significantly lower than that of Dyval2 [13]. This substantial reduction in overlap rate signifies that our framework excels at incorporating new knowledge into the generated datasets.
We also analyze the semantic embedding of the generated dataset and the original dataset, which is shown in Appendix E.
Table 4: Effectiveness of each module in UniGen.
@@
4.3 Effectiveness of Modules in UniGen
Figure 5: The percentage of different epoch counts in four datasets.
In this section, we validate the effectiveness of modules in UniGen. To simplify the analysis, our evaluation is based on the GPT-4 generated data: (1) Diversity Setting. As demonstrated in Table 4, the UniGen modules significantly enhance the diversity of the generated data. Specifically, the remote-clique score of the initially generated data stands at 0.695. However, the introduction of attribute-guided generation elevates the remote-clique score to 0.735. Furthermore, the implementation of group checking further increases this score to 0.743. (2) Overall Quality Assessment and Enhancement. To evaluate the effectiveness of our quality assessment and enhancement module, we conducted human evaluations focusing on two key aspects: (I) Comparing the quality between original and enhanced data items; (II) Assessing the reasonableness of the reflections. As illustrated in Figure 3, the results indicate that almost all reflections were deemed reasonable by the evaluators. Furthermore, over 80% of the enhanced data items were rated as superior in both datasets. These findings underscore the effectiveness of our module. (3) Difficulty Enhancement. As demonstrated in Table 6, it is observable that the performance of most of the LLMs declined when compared to their performance on the baseline-generated datasets after the application of difficulty enhancement. This result underscores the effectiveness of difficulty enhancement, which suggests its potential utility in preventing data contamination [77, 78, 79]. Such techniques may thus contribute significantly to improving the robustness of LLMs against overfitting to training datasets. (4) Code-Based Mathematical Evaluation. As depicted in Table 4, our code-based evaluation methodology has significantly enhanced the correctness of the generated data, improving from an initial accuracy of 44% to 92%. (5) Truthfulness Validation by RAG. As detailed in Table 4, the RAG-based validation corrected 4.2% of the examples, demonstrating its effectiveness. This percentage also highlights the high quality of the dataset generated by GPT-4, which contains only a few errors. The correctness of (4) and (5) are also manually evaluated, which of the details can be found in Appendix C.
In Appendix E, we explore the influence of temperature settings on data diversity and assess how effectively LLMs in UniGen adhere to user constraints. Our findings indicate that adjusting the temperature setting enhances the diversity of generated data. Additionally, LLMs within UniGen successfully follow user-imposed constraints in both single and combined scenarios.
@@ 4.4 Human Performance on Generated Dataset
As depicted in Figure 6, the performance comparison between humans and LLMs reveals distinct outcomes across various datasets. In the HellaSwag dataset, human performance slightly surpasses that of LLMs. However, in the other three datasets, LLMs demonstrate superior performance. Notably, in the GSM8K dataset, the accuracy of human responses is lower than that of the best-performing LLM. For the TruthfulQA and MMLU datasets, which require extensive knowledge, humans perform significantly worse than LLMs, which benefit from training on large, diverse corpora. More details about evaluating human performance are shown in Appendix C.
@@ 4.5 Error Analysis
To examine the errors present in the generated dataset, we conducted a human evaluation for error analysis. We observed significant factuality errors in datasets such as GSM8K, TruthfulQA, and MMLU, primarily because these datasets contain responses that are fact-based (e.g., arithmetic question answers). This observation underscores the necessity for enhancements in the accuracy of provided answers. Despite the robust instruction-following capabilities of GPT-4, it occasionally struggles with data formatting issues. Such errors could be mitigated through clearer prompts or by employing an integrated framework like LangChain***https://github.com/langchain-ai/langchain. Additionally, our analysis of the HellaSwag dataset revealed the presence of multiple viable answers for certain prompts, highlighting the need for a more comprehensive answer validation mechanism. We have discussed the potential improvement by mitigating these errors in Appendix A.
@@
4.6 Cost Analysis
We conducted a cost analysis of UniGen. Specifically, we calculated the total token usage and the corresponding cost for generating data across four datasets: MMLU, HellaSwag, TruthfulQA, and GSM8K. The details are presented in Figure 7. For a generated item without RAG-based validation and code-based evaluation, the cost is at most \$0.038 using the GPT-4-Turbo API. When incorporating RAG-based validation, the average cost per generated item increases to \$0.190, due to the large volume of tokens processed from the retrieved content. Adding code-based evaluation raises the cost to \$0.040. Overall, the total cost for generating each item, including all validation and evaluation processes, will not exceed \$0.200. This cost, although significant, is substantially lower than the cost of human labor.
Figure 7: Cost (dollar) on different epoch numbers of overall quality assessment and enhancement (Left), and the token number cost of each part in UniGen.
@@ 4.7 Application-I: Benchmarking LLMs
We present the benchmarking results based on GPT-4 and Claude3 generated data for seven popular LLMs in subsection 4.7 (the benchmarking results based on Llama3-70b’s generation are shown in Appendix E). The analysis yields several key observations:
  • $\triangleright$
    Performance decline on generated GSM8K dataset: Almost all LLMs exhibit a performance drop on the generated GSM8K dataset compared to the original. This suggests that the reasoning capabilities of many LLMs may be overstated, aligning with recent findings [80], which indicate overfitting on the GSM8K dataset by some LLMs.
  • $\triangleright$
@@
  • $\triangleright$
    Challenging nature of Claude3-generated dataset: LLMs generally perform worse on datasets generated by Claude3 compared to those by GPT-4. This may imply that some LLMs might have been trained or augmented with GPT-4 generated data (e.g., Phi-3 [81]), highlighting the unique challenge of Claude3-generated content.
@@ 4.8 Application-II: Data Augmentation
Using data augmentation with LLMs has been widely explored in previous studies [51, 82, 83]. In this section, we implement our UniGen to augment data in ten popular datasets (the details of datasets are shown in Appendix B). We include the experiment setting in Appendix D. The results of data augmentation are shown in Figure 8. We observe the following:
  • $\triangleright$
    The data augmentation powered by UniGen is effective. Performance across all ten datasets improved when trained with the UniGen–generated dataset, underscoring the efficacy of our generated data and suggesting broader potential applications for UniGen across extensive datasets.
  • $\triangleright$
    UniGen enhances LLMs from various capability aspects. The enhancements in various aspects of LLM capabilities due to the generated data are notable. For example, performance improvements in the Metatool dataset [9] (i.e., tool selection ability) indicate that UniGen can enhance agent-oriented capabilities of LLMs. Additionally, enhancements in reasoning abilities are evident in datasets such as GSM8K [19] and both the BBH (bool/casual) [84].
  • $\triangleright$
    Improvement on knowledge-intensive datasets still leaves much to be desired. The gains in datasets requiring extensive knowledge (e.g., TruthfulQA [72]) are comparatively modest. This limited improvement may be due to LLMs acquiring most of their knowledge during pretraining, and the additional 200 training samples may not significantly impact performance on related tasks. Notably, the Llama2-7b model showed a performance decline on TruthfulQA after fine-tuning, possibly due to hallucinations introduced when new knowledge is acquired during fine-tuning rather than pretraining [85]. We discussed the potential measurement for enhancing in Appendix A.

### v5

@@ 2.4 Post-Processing
Difficulty Enhancement. Given that the dataset is produced by LLMs, the complexity of the generated data is occasionally insufficient to challenge LLMs as their capabilities evolve. To address this, and inspired by prior research (wang2024benchmark; zhu2024dyval2), we implement several strategies to increase the data’s difficulty. These strategies are designed to elevate the challenges faced by LLMs in processing and responding to the data. The applied policies include: (1) Paraphrasing Question: Reformulate the phrasing to express the same idea with greater sophistication. (2) Adding Extra Context into Question: Integrate additional context or details that, while not directly aiding in the question’s resolution, enhance the question’s complexity. (3) Paraphrasing The Choices: Each option should be rephrased to reflect the same concept or idea as the original. The essence and meaning must be preserved. If an option cannot be paraphrased without altering its meaning, it should remain unchanged. (4) Adding A New Choice: Introduce a plausible but incorrect option to the existing choices to create ambiguity and require deeper understanding.
Group Checking. To mitigate the issue of high similarity among generated data items, a group-checking mechanism is implemented to identify and eliminate duplicates. Specifically, we utilize OpenAI’s text-embedding-ada-002 (text-embedding-ada-002) to compute embeddings for all generated items. Let $\mathcal{X}=\{x_{1},x_{2},\dots,x_{n}\}$ be the set of generated data items, and $\mathbf{e}_{i}$ be the embedding of item $x_{i}$ computed via text-embedding-ada-002. We define the similarity matrix $\mathbf{S}$ where the element $s_{ij}$ is given by $s_{ij}=\sqrt{\sum_{k=1}^{d}(e_{ik}-e_{jk})^{2}}$, representing the Euclidean distance between the embeddings of items $x_{i}$ and $x_{j}$. Data items exhibiting a similarity exceeding a predefined threshold $\theta$ are filtered out to ensure diversity within the dataset. Formally, if $s_{ij}<\theta$ for any pair $(i,j)$, at least one of the items $x_{i}$ or $x_{j}$ is randomly removed from the final dataset.
3 Experiments and Applications
@@ 3.1 Experimental Setup
To thoroughly evaluate the effectiveness of DataGen, we carefully select four representative benchmark datasets: GSM8K (cobbe2021gsm8k), TruthfulQA (lin2022truthfulqa), MMLU (hendryckstest2021), and HellaSwag (zellers2019hellaswag). Each dataset uniquely contributes to language model assessment, covering dimensions from mathematical problem-solving and factual accuracy verification to extensive language understanding and commonsense reasoning. We show the details of these four datasets in Appendix C. For dataset generation, we utilize GPT-4 (ChatGPT), Claude3-Opus (Claude), and Llama3-70b (LLAMA3), as these LLMs are among the most robust available, exhibiting exceptional ability to follow instructions. For benchmarking, our study utilizes eight popular models from notable entities in the AI domain (the details are shown in Appendix C.), reflecting a mix of open-source and proprietary technologies. The number of generated data items and more details are shown in Appendix E. Note that difficulty enhancement is not applied to the generated data for benchmarking. We will discuss the effectiveness of difficult enhancement in Section 3.3. All LLMs utilized for generation share the same prompt templates.
@@ 3.2 Characterizing Generated Data
Length. As depicted in 4(a), the length distribution of all generated datasets approximates a normal distribution. Notably, except for the HellaSwag dataset (as the length of the original HellaSwag dataset looks like a bimodal distribution), the distributions of other datasets closely resemble those of their original datasets. This similarity indicates that DataGen effectively mimics the distribution of the original data, thereby enhancing the reliability of the generated datasets.
Figure 5: Semantic embedding of different datasets. We use OpenAI’s text-embedding-ada-002 (text-embedding-ada-002) to obtain text embedding.
Diversity. Analogous to the length distribution, the distribution of the self-BLEU score (zhu2018texygen) (as depicted in 4(b))—a metric employed to assess text diversity—indicates that the diversity of the generated data closely aligns with that of the original dataset. This alignment underscores the exceptional capability of DataGen to replicate the diversity inherent in the original dataset, demonstrating its effectiveness in producing varied textual content. Additionally, we utilize the remote-clique metric, as applied in prior research (cevallos2018diversity), to measure the diversity of the generated data. The related statistics are presented in Table 2. Observations reveal that the remote-clique scores of the original and generated data are closely matched, with less than 10% variance, affirming that our generated data maintains a high level of diversity comparable to the original dataset.
Knowledge Richness Introduced. In contrast to prior research (zhu2024dyval; zhu2024dyval2; wang2024benchmark), DataGen innovates by generating entirely new data items, rather than merely modifying existing answers. This approach introduces novel scenarios and knowledge. We assess the knowledge richness of the data generated by DataGen and compared it to the previous study (i.e., Dyval2 (zhu2024dyval2)) by calculating the entity overlap rate—how many entities appear both in the generated and original data. A lower overlap rate indicates that the framework is introducing more new knowledge. According to our findings, presented in Table 3, DataGen demonstrates an average overlap rate of only 3.83%, significantly lower than that of Dyval2 (zhu2024dyval2). This substantial reduction in overlap rate signifies that our framework excels at incorporating new knowledge into the generated datasets.
Influence of temperature. We examine the impact of temperature settings on the diversity of data generated by GPT-4. For this purpose, we select a few items from the TruthfulQA dataset to use as examples in few-shot learning. We conduct experiments using temperature settings of 0 and 1. Our findings indicate that the Remote-Clique score (cevallos2018diversity) at a temperature of 0 is 0.683, whereas, at a temperature of 1, it increases to 0.721. This suggests that adjusting the temperature setting can significantly enhance the diversity of the generated data.
Table 3: The knowledge richness comparison between different principles in DyVal 2 (zhu2024dyval2) and DataGen. The principle 1, 2, 3, and 4 are paraphrasing questions, paraphrasing choices, adding extra context to questions, and adding a new choice.
@@
DataGen 5.40% 3.30% 2.80% 3.83%
3.3 Effectiveness of Modules in DataGen
Figure 6: The percentage of different epoch counts in four datasets.
In this section, we validate the effectiveness of modules in DataGen. To simplify the analysis, our evaluation is based on the GPT-4 generated data: (1) Diversity Setting. As demonstrated in Table 4, the DataGen modules significantly enhance the diversity of the generated data. Specifically, the remote-clique score of the initially generated data stands at 0.695. However, the introduction of attribute-guided generation elevates the remote-clique score to 0.735. Furthermore, the implementation of group checking further increases this score to 0.743. (2) Overall Quality Assessment and Enhancement. To evaluate the effectiveness of our quality assessment and enhancement module, we conducted human evaluations focusing on two key aspects: (I) Comparing the quality between original and enhanced data items; (II) Assessing the reasonableness of the reflections. As illustrated in Figure 3, the results indicate that almost all reflections were deemed reasonable by the evaluators. Furthermore, over 80% of the enhanced data items were rated as superior in both datasets. These findings underscore the effectiveness of our module. (3) Difficulty Enhancement. As demonstrated in Table 6, it is observable that the performance of most of the LLMs declined when compared to their performance on the baseline-generated datasets after the application of difficulty enhancement. This result underscores the effectiveness of difficulty enhancement, which suggests its potential utility in preventing data contamination (dong2024generalization; golchin2024time; xu2024benchmarking). Such techniques may thus contribute significantly to improving the robustness of LLMs against overfitting to training datasets. (4) Code-Based Mathematical Evaluation. As depicted in Table 4, our code-based evaluation methodology has significantly enhanced the correctness of the generated data, improving from an initial accuracy of 44% to 92%. (5) Truthfulness Validation by RAG. As detailed in Table 4, the RAG-based validation corrected 4.2% of the examples, demonstrating its effectiveness. This percentage also highlights the high quality of the dataset generated by GPT-4, which contains only a few errors. The correctness of (4) and (5) are also manually evaluated, which of the details can be found in Appendix D.
In Appendix F, we investigate the impact of temperature settings on data diversity and evaluate the adherence of LLMs in DataGen to user constraints. Our findings reveal that adjusting the temperature setting enhances the diversity of generated data. Furthermore, LLMs within DataGen effectively follow user-imposed constraints in both individual and combined scenarios. We also provide a cost analysis of DataGen in Appendix F, demonstrating that DataGen generates datasets at a significantly low cost.
Table 4: Effectiveness of each module in DataGen.
@@ 3.4 Human Performance on Generated Dataset
As depicted in Figure 7, the performance comparison between humans and LLMs reveals distinct outcomes across various datasets. In the HellaSwag dataset, human performance slightly surpasses that of LLMs. However, in the other three datasets, LLMs demonstrate superior performance. Notably, in the GSM8K dataset, the accuracy of human responses is lower than that of the best-performing LLM. For the TruthfulQA and MMLU datasets, which require extensive knowledge, humans perform significantly worse than LLMs, which benefit from training on large, diverse corpora. More details about evaluating human performance are shown in Appendix D.
@@ 3.5 Error Analysis
To examine the errors present in the generated dataset, we conducted a human evaluation for error analysis. We observe significant factuality errors in datasets such as GSM8K, TruthfulQA, and MMLU, primarily because these datasets contain responses that are fact-based (e.g., arithmetic question answers). This observation underscores the necessity for enhancements in the accuracy of provided answers. Despite the robust instruction-following capabilities of GPT-4, it occasionally struggles with data formatting issues. Such errors could be mitigated through clearer prompts or by employing an integrated framework like LangChain***https://github.com/langchain-ai/langchain. Additionally, our analysis of the HellaSwag dataset revealed the presence of multiple viable answers for certain prompts, highlighting the need for a more comprehensive answer validation mechanism. We discuss the potential improvement by mitigating these errors in Appendix A.
3.6 Cost Ablation Analysis
We conduct a cost analysis of DataGen. Specifically, we calculate the total token usage and the corresponding cost for generating data across four datasets: MMLU, HellaSwag, TruthfulQA, and GSM8K. The details are presented in Figure 8.
For a generated item without RAG-based validation and code-based evaluation, the cost is at most \$0.038 using the GPT-4-Turbo API. When incorporating RAG-based validation, the average cost per generated item increases to \$0.190, due to the large volume of tokens processed from the retrieved content. Adding code-based evaluation raises the cost to \$0.040. Overall, the total cost for generating each item, including all validation and evaluation processes, will not exceed \$0.200. This cost, although significant, is substantially lower than the cost of human labor.
Figure 8: Cost (dollar) on different epoch numbers of overall quality assessment and enhancement (Left), and the token number cost of each part in DataGen.
@@ 3.7 Application-I: Benchmarking LLMs
We present the benchmarking results based on GPT-4 and Claude3 generated data for seven popular LLMs in Table 7 (the benchmarking results based on Llama3-70b’s generation are shown in Appendix F). The analysis yields several key observations:
  • Performance decline on generated GSM8K dataset: Almost all LLMs exhibit a performance drop on the generated GSM8K dataset compared to the original. This suggests that the reasoning capabilities of many LLMs may be overstated, aligning with recent findings (zhang2024careful; mirzadeh2024gsm; zhang2024careful), which indicate overfitting on the GSM8K dataset by some LLMs.
@@
  • Challenging nature of Claude3-generated dataset: LLMs generally perform worse on datasets generated by Claude3 compared to those by GPT-4. This may imply that some LLMs might have been trained or augmented with GPT-4 generated data (e.g., Phi-3 (abdin2024phi3)), highlighting the unique challenge of Claude3-generated content.
@@ 3.8 Application-II: Data Augmentation
Using data augmentation with LLMs has been widely explored in previous studies (dai2023auggpt; whitehouse2023llmpowered; moller-etal-2024-parrot). In this section, we implement our DataGen to augment data in ten popular datasets (the details of datasets are shown in Appendix C). We include the experiment setting in Appendix E. From Figure 9, we can observe that:
  • The data augmentation powered by DataGen is effective. Performance across all ten datasets improved when trained with the DataGen-generated dataset, highlighting the efficacy of our generated data and indicating broader potential applications for DataGen across extensive datasets.
  • DataGen enhances LLMs from various capability aspects. The enhancements in various aspects of LLM capabilities due to the generated data are notable. For example, performance improvements in the Metatool dataset (huang2024metatool) (i.e., tool selection ability) indicate that DataGen can enhance agent-oriented capabilities of LLMs. Additionally, enhancements in reasoning abilities are evident in datasets such as GSM8K (cobbe2021gsm8k) and both the BBH (bool/casual) (suzgun2022challenging).
  • Improvement on knowledge-intensive datasets still leaves much to be desired. The gains in datasets requiring extensive knowledge (e.g., TruthfulQA (lin2022truthfulqa)) are comparatively modest. This limited improvement may be due to LLMs acquiring most of their knowledge during pretraining, and the additional 200 training samples may not significantly impact performance on related tasks. Notably, the Llama2-7b model shows a performance decline on TruthfulQA after fine-tuning, possibly due to hallucinations introduced when new knowledge is acquired during fine-tuning rather than pretraining (gekhman2024does). We discuss the potential measurement for enhancing in Appendix A.
Moreover, we extend our analysis to include general instruction tuning data. Specifically, we utilize the alpaca dataset alpaca for additional fine-tuning on the Llama3-base model and evaluated the outcomes using the MT-Bench zheng2023judging. The “genset” model, fine-tuned on 1,000 data points generated by DataGen, consistently outperforms the “original” model, which is fine-tuned on an equivalent sample of 1,000 existing data points from the alpaca dataset. This comparison demonstrates that our framework effectively generates high-quality, diverse instruction-tuning data, demonstrating its practical utility in enhancing model performance.
Table 8: Model performance scores in MTBench.
Model First Turn Score Second Turn Score Average Score
llama3-7b-base 2.325 1.744 2.038
llama3-7b-alpaca-original 6.049 4.450 5.259
llama3-7b-alpaca-genset 6.981 5.825 6.403